# OpenScan: A Benchmark for Generalized Open-Vocabulary 3D Scene Understanding

### v2

@@ Appendix B Implementation Details
| Method | Affordance | Property | Type | Manner | Synonym | Requirement | Element | Mean | ||||||||||||||||
| AP | AP50 | AP25 | AP | AP50 | AP25 | AP | AP50 | AP25 | AP | AP50 | AP25 | AP | AP50 | AP25 | AP | AP50 | AP25 | AP | AP50 | AP25 | AP | AP50 | AP25 | |
| OpenMask3D | 7.2 | 9.1 | 10.4 | 7.5 | 10.0 | 11.6 | 8.5 | 11.2 | 13.0 | 12.8 | 15.4 | 17.4 | 16.9 | 19.7 | 20.6 | 13.0 | 16.0 | 18.9 | 12.2 | 15.4 | 17.1 | 9.7 | 12.2 | 13.9 |
| Upper Bound | 19.3 | 24.3 | 27.6 | 26.7 | 34.2 | 37.8 | 18.5 | 24.3 | 27.4 | 27.7 | 35.1 | 38.6 | 29.4 | 34.7 | 37.4 | 26.5 | 34.4 | 39.4 | 22.1 | 27.2 | 31.2 | 21.6 | 27.4 | 30.9 |
| 12.1 | 15.2 | 17.2 | 19.2 | 24.2 | 26.2 | 10.0 | 13.1 | 14.4 | 14.9 | 19.7 | 21.2 | 12.5 | 15.0 | 16.8 | 13.5 | 18.4 | 20.5 | 9.9 | 11.8 | 14.1 | 11.7 | 14.9 | 16.7 | |
| SAI3D | 5.3 | 8.4 | 10.5 | 5.8 | 8.3 | 10.7 | 7.8 | 11.4 | 13.4 | 11.3 | 15.7 | 18.2 | 10.0 | 16.7 | 20.0 | 10.0 | 15.3 | 18.7 | 8.7 | 13.6 | 16.0 | 7.6 | 11.5 | 13.8 |
| Upper Bound | 14.6 | 22.0 | 28.3 | 17.7 | 27.7 | 35.4 | 17.4 | 26.0 | 33.2 | 18.9 | 28.4 | 36.0 | 19.3 | 28.3 | 32.9 | 19.2 | 30.1 | 39.5 | 16.8 | 25.1 | 32.4 | 16.8 | 25.4 | 32.4 |
| 9.3 | 13.6 | 17.8 | 11.9 | 19.4 | 24.7 | 9.6 | 14.6 | 19.8 | 7.6 | 12.7 | 17.8 | 9.3 | 11.6 | 12.9 | 9.2 | 14.8 | 20.8 | 8.1 | 11.5 | 16.4 | 9.2 | 13.9 | 18.6 | |
| MaskClustering | 6.2 | 10.7 | 13.7 | 7.0 | 12.3 | 15.8 | 7.1 | 13.3 | 17.7 | 11.1 | 18.4 | 23.1 | 16.2 | 30.3 | 36.6 | 11.3 | 21.8 | 28.2 | 7.4 | 13.5 | 17.2 | 7.9 | 14.4 | 18.5 |
| Upper Bound | 15.4 | 28.6 | 37.2 | 21.1 | 38.5 | 48.8 | 13.9 | 26.8 | 35.0 | 18.0 | 33.5 | 44.8 | 18.0 | 34.2 | 45.0 | 26.3 | 49.7 | 61.6 | 17.6 | 32.5 | 42.7 | 16.9 | 31.6 | 41.0 |
| 9.2 | 17.9 | 23.5 | 14.1 | 26.2 | 33.0 | 6.8 | 13.5 | 17.3 | 6.9 | 15.1 | 21.7 | 1.8 | 3.9 | 8.4 | 15.0 | 27.9 | 33.4 | 10.2 | 19.0 | 25.5 | 9.0 | 17.2 | 22.5 | |
| Open3DIS | 11.9 | 14.8 | 16.7 | 12.8 | 16.0 | 16.8 | 14.2 | 17.9 | 20.2 | 19.2 | 22.3 | 24.2 | 26.7 | 30.6 | 33.1 | 19.2 | 24.1 | 25.5 | 18.7 | 21.9 | 24.7 | 15.4 | 18.9 | 20.9 |
| Upper Bound | 23.9 | 29.9 | 32.7 | 36.0 | 43.8 | 47.1 | 24.5 | 30.6 | 34.5 | 34.9 | 43.1 | 46.0 | 36.1 | 42.6 | 44.9 | 35.0 | 43.8 | 47.2 | 28.0 | 33.5 | 38.6 | 27.7 | 34.1 | 37.6 |
| 12.0 | 15.1 | 16.0 | 23.2 | 27.8 | 30.3 | 10.3 | 12.7 | 14.3 | 15.7 | 20.8 | 21.8 | 9.4 | 12.0 | 11.8 | 15.8 | 19.7 | 21.7 | 9.3 | 11.6 | 13.9 | 12.3 | 15.2 | 16.7 | |
Appendix C Additional Experimental Results
C.1 Results of Radar Charts.
Figure 8 presents radar charts of the main results on our OpenScan benchmark. The charts compare performance across AP, AP50, and AP25 for OpenMask3D [34], SAI3D [41], MaskClustering [36], and Open3DIS [24]. Our results demonstrate that Open3DIS achieves the strongest performance across eight linguistic aspects, particularly in terms of AP and AP50. Meanwhile, MaskClustering exhibits competitive performance in AP50 and AP25, with notable strengths in the synonym and requirement aspects.
@@ Appendix C Additional Experimental Results
In this paper, we adopt query templates (e.g., “this term is made of wood”) as the default experimental configuration. For template-free evaluation on the GOV-3D benchmark (e.g., “wood”), detailed results are provided in Table 7. We evaluate OpenMask3D [34], SAI3D [41], MaskClustering [36], and Open3DIS [24] for 3D instance segmentation. Our experiments show that Open3DIS achieves the highest AP, AP50, and AP25 scores across every linguistic aspect. This performance aligns with its strong performance in the GOV-3D task with query templates.
@@ C.3 Results of Visual Attributes
We present comparative visual attribute results for the material aspect on our OpenScan benchmark in Table 8. Our evaluation involves OpenMask3D [34], SAI3D [41], MaskClustering [36], and Open3DIS [24] across 10 material attributes. It demonstrates that these OV-3D models perform strongly on the “porcelain” material, indicating that the visual information of the “porcelain” material in 3D objects (e.g., “toilet” and “bathtub”) is more distinguishable than that of other materials. However, these OV-3D models struggle to accurately segment the “stone” material. This difficulty stems from the fact that stone is commonly associated with large 3D regions (e.g., “wall” and “floor”), which are often neglected following the common practice [31, 34, 41] during 3D segmentation. These OV-3D models cannot correctly segment these large 3D areas, resulting in low results of the “stone” material. Notably, Open3DIS shows impressive results on each material compared to other OV-3D models, aligning with its strong performance in the classic OV-3D task.
@@ C.4 Results of Upper Bound
During the 3D prediction in the GOV-3D task, we query the attributes to obtain the attribute-related 3D mask predictions. For annotating the OpenScan benchmark, object classes are associated with corresponding attributes using the ConceptNet [32] database. Conversely, each attribute query can also be associated with the corresponding object classes. Therefore, we can replace the attribute queries with the ground truth attribute-related object classes from ConceptNet to finalize 3D mask results. We exclude the material aspect since the related object classes of material in the ConceptNet database are limited. We serve this setting as our upper bound performance. Table 9 shows the comparison of baseline methods OpenMask3D [34], SAI3D [41], MaskClustering [36], and Open3DIS [24] with their upper bounds, highlighting significant performance gaps that underscore the potential for attribute-aware 3D reasoning.
Appendix D Additional Benchmark Details
@@ D.1 Does OpenScan Represent 200 Object Classes From ScanNet200 Well Enough?
During the annotation of our OpenScan benchmark, object classes from ScanNet200 [30] are labeled with attributes using the ConceptNet [32] database or through manual annotation. Figure 9 shows the number of attributes per object class from ScanNet200 in our OpenScan benchmark. Notably, most object classes from ScanNet200 are annotated with more than one attribute in our OpenScan, indicating that our OpenScan benchmark adequately represents object classes from ScanNet200. Besides, the object class has up to seven attributes (i.e., “bicycle”, and “ball”) in our OpenScan benchmark.
@@
Figure 10 summarizes the distribution of attributes per object in our OpenScan benchmark. Most objects are annotated with 1-6 attributes.
@@
Figure 11 presents the distribution of attributes per scene in our OpenScan benchmark. It demonstrates that the attributes in 3D scenes are semantically rich.
D.4 Benchmark Formats
Figure 13 shows an example of our OpenScan benchmark formats. Our OpenScan is formatted in the JSON file. Each target 3D object is annotated with the following items:
@@ D.4 Benchmark Formats
D.5 Additional Benchmark Samples
D.6 Benchmark Details
We construct our OpenScan benchmark based on ScanNet200 [30] across eight linguistic aspects. We present all attributes and their corresponding query templates in our OpenScan benchmark: Table 10 displays the affordance and property aspects; Table 11 shows the type, manner, and synonym aspects; and Table 12 presents the requirement, element, and material aspects. The “object” in the query template is replaced with “this term” in our experiments.
Appendix E Limitations and Future Work
Our benchmark is currently constructed solely on the ScanNet200 benchmark with limited 3D indoor scene. It would be beneficial to increase the scale of our benchmark to include a wider variety of 3D scenes and objects. In future work, we plan to extend our OpenScan benchmark to encompass more diverse scenes by incorporating indoor 3D datasets such as ScanNet++ [40] and Matterport3D [5]. Our mature annotation procedures can be readily adapted to these datasets. Moreover, we aim to evaluate current OV-3D models on our GOV-3D task, particularly examining performance variations when using higher point resolutions in ScanNet++ [40] and larger scene areas in Matterport3D [5].
Appendix F Broader Impact

### v3

@@ C Additional Experimental Results
| Method | Affordance | Property | Type | Manner | Synonym | Requirement | Element | Mean |
| UpB ✗ / ✓ | UpB ✗ / ✓ | UpB ✗ / ✓ | UpB ✗ / ✓ | UpB ✗ / ✓ | UpB ✗ / ✓ | UpB ✗ / ✓ | UpB ✗ / ✓ | |
| AP | ||||||||
| OpenMask3D | 7.2 / 19.3 | 7.5 / 26.7 | 8.5 / 18.5 | 12.8 / 27.7 | 16.9 / 29.4 | 13.0 / 26.5 | 12.2 / 22.1 | 9.7 / 21.6 |
| SAI3D | 5.3 / 14.6 | 5.8 / 17.7 | 7.8 / 17.4 | 11.3 / 18.9 | 10.0 / 19.3 | 10.0 / 19.2 | 8.7 / 16.8 | 7.6 / 16.8 |
| MaskClustering | 6.2 / 15.4 | 7.0 / 21.1 | 7.1 / 13.9 | 11.1 / 18.0 | 16.2 / 18.0 | 11.3 / 26.3 | 7.4 / 17.6 | 7.9 / 16.9 |
| Open3DIS | 11.9 / 23.9 | 12.8 / 36.0 | 14.2 / 24.5 | 19.2 / 34.9 | 26.7 / 36.1 | 19.2 / 35.0 | 18.7 / 28.0 | 15.4 / 27.7 |
| AP50 | ||||||||
| OpenMask3D | 9.1 / 24.3 | 10.0 / 34.2 | 11.2 / 24.3 | 15.4 / 35.1 | 19.7 / 34.7 | 16.0 / 34.4 | 15.4 / 27.2 | 12.2 / 27.4 |
| SAI3D | 8.4 / 22.0 | 8.3 / 27.7 | 11.4 / 26.0 | 15.7 / 28.4 | 16.7 / 28.3 | 15.3 / 30.1 | 13.6 / 25.1 | 11.5 / 25.4 |
| MaskClustering | 10.7 / 28.6 | 12.3 / 38.5 | 13.3 / 26.8 | 18.4 / 33.5 | 30.3 / 34.2 | 21.8 / 49.7 | 13.5 / 32.5 | 14.4 / 31.6 |
| Open3DIS | 14.8 / 29.9 | 16.0 / 43.8 | 17.9 / 30.6 | 22.3 / 43.1 | 30.6 / 42.6 | 24.1 / 43.8 | 21.9 / 33.5 | 18.9 / 34.1 |
| AP25 | ||||||||
| OpenMask3D | 10.4 / 27.6 | 11.6 / 37.8 | 13.0 / 27.4 | 17.4 / 38.6 | 20.6 / 37.4 | 18.9 / 39.4 | 17.1 / 31.2 | 13.9 / 30.9 |
| SAI3D | 10.5 / 28.3 | 10.7 / 35.4 | 13.4 / 33.2 | 18.2 / 36.0 | 20.0 / 32.9 | 18.7 / 39.5 | 16.0 / 32.4 | 13.8 / 32.4 |
| MaskClustering | 13.7 / 37.2 | 15.8 / 48.8 | 17.7 / 35.0 | 23.1 / 44.8 | 36.6 / 45.0 | 28.2 / 61.6 | 17.2 / 42.7 | 18.5 / 41.0 |
| Open3DIS | 16.7 / 32.7 | 16.8 / 47.1 | 20.2 / 34.5 | 24.2 / 46.0 | 33.1 / 44.9 | 25.5 / 47.2 | 24.7 / 38.6 | 20.9 / 37.6 |
| LLM ✗ / ✓ | LLM ✗ / ✓ | LLM ✗ / ✓ | LLM ✗ / ✓ | LLM ✗ / ✓ | LLM ✗ / ✓ | LLM ✗ / ✓ | LLM ✗ / ✓ | LLM ✗ / ✓ | |
| AP | |||||||||
| OpenMask3D | 7.2 / 10.8 | 7.5 / 13.9 | 8.5 / 8.5 | 12.8 / 14.1 | 16.9 / 25.7 | 13.0 / 13.4 | 12.2 / 12.1 | 18.8 / 10.7 | 9.9 / 11.7 |
| SAI3D | 5.3 / 7.2 | 5.8 / 9.9 | 7.8 / 7.5 | 11.3 / 14.5 | 10.0 / 16.1 | 10.0 / 9.3 | 8.7 / 7.9 | 11.3 / 6.6 | 7.7 / 8.6 |
| MaskClustering | 6.2 / 8.2 | 7.0 / 8.5 | 7.1 / 7.9 | 11.1 / 9.1 | 16.2 / 11.1 | 11.3 / 16.1 | 7.4 / 11.6 | 12.1 / 8.9 | 8.1 / 9.5 |
| Open3DIS | 11.9 / 15.1 | 12.8 / 22.7 | 14.2 / 15.4 | 19.2 / 21.6 | 26.7 / 31.0 | 19.2 / 21.3 | 18.7 / 17.6 | 28.3 / 18.7 | 15.8 / 17.8 |
| AP50 | |||||||||
| OpenMask3D | 9.1 / 13.5 | 10.0 / 18.7 | 11.2 / 11.0 | 15.4 / 17.9 | 19.7 / 30.7 | 16.0 / 17.3 | 15.4 / 15.0 | 22.1 / 12.6 | 12.5 / 14.7 |
| SAI3D | 8.4 / 10.4 | 8.3 / 15.7 | 11.4 / 10.9 | 15.7 / 20.7 | 16.7 / 24.0 | 15.3 / 15.1 | 13.6 / 12.6 | 17.1 / 10.0 | 11.6 / 12.8 |
| MaskClustering | 10.7 / 14.3 | 12.3 / 16.3 | 13.3 / 15.1 | 18.4 / 16.1 | 30.3 / 21.9 | 21.8 / 30.5 | 13.5 / 21.2 | 20.6 / 15.7 | 14.6 / 17.5 |
| Open3DIS | 14.8 / 18.7 | 16.0 / 28.6 | 17.9 / 18.8 | 22.3 / 26.2 | 30.6 / 36.9 | 24.1 / 26.7 | 21.9 / 21.0 | 33.6 / 21.8 | 19.3 / 21.7 |
| AP25 | |||||||||
| OpenMask3D | 10.4 / 15.3 | 11.6 / 21.1 | 13.0 / 14.3 | 17.4 / 19.9 | 20.6 / 33.2 | 18.9 / 20.3 | 17.1 / 17.0 | 25.0 / 14.1 | 14.2 / 17.1 |
| SAI3D | 10.5 / 13.2 | 10.7 / 20.2 | 13.4 / 14.6 | 18.2 / 23.3 | 20.0 / 28.3 | 18.7 / 18.9 | 16.0 / 15.8 | 22.9 / 13.4 | 14.1 / 16.2 |
| MaskClustering | 13.7 / 17.7 | 15.8 / 20.4 | 17.7 / 18.5 | 23.1 / 22.4 | 36.6 / 26.7 | 28.2 / 36.0 | 17.2 / 25.5 | 25.6 / 19.9 | 18.7 / 21.5 |
| Open3DIS | 16.7 / 20.4 | 16.8 / 30.2 | 20.2 / 20.4 | 24.2 / 28.2 | 33.1 / 39.2 | 25.5 / 28.5 | 24.7 / 23.4 | 36.7 / 24.3 | 21.4 / 23.6 |
@@
In this paper, we adopt query templates (e.g., “this term is made of wood”) as the default experimental configuration. For template-free evaluation on the GOV-3D benchmark (e.g., “wood”), detailed results are provided in Table A. We evaluate OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024) for 3D instance segmentation. Our experiments show that Open3DIS achieves the highest AP, AP50, and AP25 scores across every linguistic aspect. This performance aligns with its strong performance in the GOV-3D task with query templates.
@@ C.3 Results of Visual Attributes
We present comparative visual attribute results for the material aspect on our OpenScan benchmark in Table B. Our evaluation involves OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024) across 10 material attributes. It demonstrates that these OV-3D models perform strongly on the “porcelain” material, indicating that the visual information of the “porcelain” material in 3D objects (e.g., “toilet” and “bathtub”) is more distinguishable than that of other materials. However, these OV-3D models struggle to accurately segment the “stone” material. This difficulty stems from the fact that stone is commonly associated with large 3D regions (e.g., “wall” and “floor”), which are often neglected following the common practice (Schult et al. 2023; Takmaz et al. 2023; Yin et al. 2024) during 3D segmentation. These OV-3D models cannot correctly segment these large 3D areas, resulting in low results of the “stone” material. Notably, Open3DIS shows impressive results on each material compared to other OV-3D models, aligning with its strong performance in the classic OV-3D task.
@@ C.4 Results of Upper Bound
During the 3D prediction in the GOV-3D task, we query the attributes to obtain the attribute-related 3D mask predictions. For annotating the OpenScan benchmark, object classes are associated with corresponding attributes using the ConceptNet (Speer, Chin, and Havasi 2017) database. Conversely, each attribute query can also be associated with the corresponding object classes. Therefore, we can replace the attribute queries with the ground truth attribute-related object classes from ConceptNet to finalize 3D mask results. We exclude the material aspect since the related object classes of material in the ConceptNet database are limited. We serve this setting as our upper bound performance. Table C shows the comparison of baseline methods OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024) with their upper bounds, highlighting significant performance gaps that underscore the potential for attribute-aware 3D reasoning.
C.5 Results of Introducing LLM for Attribute Understanding
In our failure case analysis in the main paper, we observe that the OV-3D model Open3DIS (Nguyen et al. 2024) can identify the object classes (e.g., “piano”) in the OV-3D task but fails to recognize the associated object attributes (e.g., “this term has 88 keys”) in the GOV-3D task. This suggests a promising direction for improving GOV-3D performance by leveraging large language models (LLMs) to perform high-level reasoning, transforming the GOV-3D attribute queries (e.g., “this term has 88 keys”) back to the OV-3D class queries (e.g., “piano”). To this end, we design an experiment where an LLM (i.e., Vicuna-7B (Chiang et al. 2023)) is prompted to map object attributes to corresponding object classes. Given an attribute [attribute], the LLM is prompted as:
Q: Given an object’s attribute [attribute], please output the related object’s classes in the indoor scene separated by commas.
The LLM will generate a list of object classes corresponding to the input attribute [attribute]. We then format the object classes into a sentence as a query for evaluation. We utilize the baseline methods OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024) to compare their performance on whether introducing LLM for attribute understanding. As shown in Table D, introducing LLM can improve the attribute understanding performance across most linguistic aspects of the GOV-3D task. However, performance declines in the material aspect, as LLMs rely exclusively on linguistic inputs and lack visual context required to differentiate material properties (e.g.“wooden chair” and “plastic chair”). Additionally, the LLM’s output can be noisy and inconsistent, occasionally producing object classes unrelated to the input attribute, which degrades performance in some linguist aspects.
C.6 Results of Weighted Mean
In the OpenScan benchmark, we observe disparities in the attribute annotations for linguistic aspects (e.g., “affordance” and “synonym”). To address the imbalance in attribute annotations, we introduce a weighted mean score (w-Mean) metric that normalizes contributions based on annotation counts. For linguist aspects with annotation counts and corresponding scores , the w-Mean is computed as:
As shown in Table E, applying the w-Mean metric improves the performance of baseline methods, including OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024). This improvement steams from the w-Mean metric’s ability to normalize the contribution of each linguistic aspect based on its annotation count, thus mitigating biases from attributes and enhancing the robustness of performance evaluation across linguistic aspects in the GOV-3D task.
| Method | Mean | w-Mean | ||||
| AP | AP50 | AP25 | AP | AP50 | AP25 | |
| OpenMask3D | 9.9 | 12.5 | 14.2 | 12.3 | 15.0 | 17.1 |
| SAI3D | 7.7 | 11.6 | 14.1 | 8.6 | 13.0 | 16.4 |
| MaskClustering | 8.1 | 14.6 | 18.7 | 9.0 | 16.0 | 20.3 |
| Open3DIS | 15.8 | 19.3 | 21.4 | 19.1 | 23.1 | 25.5 |
C.7 Additional Failure Cases Analysis
As illustrated in Figure C, the Open3DIS model(Nguyen et al. 2024) for OV-3D exhibits limitations in the GOV-3D task under specific conditions. Specifically, the model struggles to generate accurate 3D masks when: (a) the attribute query requires complex commonsense knowledge (e.g., “this term requires water and sun”), resulting in failure to predict 3D masks; (b) the target 3D object contains noisy geometry, such as 3D holes or irregular 3D structures, leading to partially incorrect 3D masks; and (c) the target object is small, providing insufficient geometric detail for segmentation, causing the model to fail in predicting 3D masks. In contrast, Open3DIS correctly predicts 3D masks for attribute-related class queries in the OV-3D task under these scenarios, underscoring the challenge of the GOV-3D task.
D Additional Benchmark Details
@@ D.1 Does OpenScan Represent 200 Object Classes From ScanNet200 Well Enough?
During the annotation of our OpenScan benchmark, object classes from ScanNet200 (Rozenberszki et al. 2022) are labeled with attributes using the ConceptNet (Speer, Chin, and Havasi 2017) database and manual annotation. Figure D shows the number of attributes per object class from ScanNet200 in our OpenScan benchmark. Notably, most object classes from ScanNet200 are annotated with more than one attribute in our OpenScan, indicating that our OpenScan benchmark adequately represents object classes from ScanNet200. Besides, the object class has up to seven attributes (i.e., “bicycle”, and “ball”) in our OpenScan benchmark.
@@
Figure E summarizes the distribution of attributes per object in our OpenScan benchmark. The majority of objects have 1–6 attributes.
@@
Figure F presents the distribution of attributes per scene in our OpenScan benchmark. It demonstrates that the attributes in 3D scenes are semantically rich.
D.4 Attribute Verification in Benchmark Annotation
During the attribute annotation process, we leverage a knowledge graph to automatically generate object-related attributes. However, the initial attribute set often contains noise, including attributes that are irrelevant, ambiguous, or semantically inconsistent with the related object classes. To address this, we conduct a meticulous manual verification process to refine the attribute set, ensuring semantic consistency and coherence in our OpenScan benchmark.
As shown in Figure G, our OpenScan benchmark generates 528 attributes initially. After attribute verification, 341 attributes are retained, resulting in an overall reduction of 35.4% of noisy attributes. Notably, the affordance aspect exhibits high noise level, with 42.5% of attributes being filtered out, suggesting that the affordance attributes are particularly prone to ambiguity due to the diverse nature of affordance candidates within the knowledge graph. In contrast, all synonym attributes are retained during verification. This robustness is attributed to the high semantic similarity between synonym attributes and their corresponding object classes, ensuring reliable alignment during the generation process.
D.5 Additional Benchmark Samples
We provide additional samples of our OpenScan benchmark. Figure H presents the examples of objects and their corresponding attributes. Figure J displays the affordance, property, type, and manner aspects, while Figure K shows the synonym, requirement, element, and material aspects.
D.6 Benchmark Formats
Figure I shows an example of our OpenScan benchmark formats. Our OpenScan is formatted in the JSON file. Each target 3D object is annotated with the following items:
@@ D.6 Benchmark Formats
D.7 Benchmark Details
We construct our OpenScan benchmark based on ScanNet200 (Rozenberszki et al. 2022) across eight linguistic aspects. We present all attributes and their corresponding query templates in our OpenScan benchmark: Table F displays the affordance and property aspects; Table G shows the type, manner, and synonym aspects; and Table H presents the requirement, element, and material aspects. The “object” in the query template is replaced with “this term” in our experiments.
E Additional Related Work
Open-Vocabulary 2D Understanding Benchmarks. Open-vocabulary 2D understanding refers to the task of detecting or segmenting novel object classes that are not present in the training dataset. For the object detection task, COCO (Lin et al. 2014) and LVIS (Gupta, Dollar, and Girshick 2019) are two widely used datasets. For the image segmentation task, popular datasets include COCO (Lin et al. 2014), ADE20k (Zhou et al. 2019), PASCAL-VOC (Everingham et al. 2015), and Cityscapes (Cordts et al. 2016). However, these benchmarks primarily evaluate the model’s open-vocabulary ability but do not explicitly assess its capability to recognize specific object characteristics. PACO (Ramanathan et al. 2023) introduces a 2D segmentation benchmark that focuses on parts and attributes of common objects. Inspired by PACO (Ramanathan et al. 2023), FG-OVD (Bianchi et al. 2024) presents a challenging task and benchmark for fine-grained open-vocabulary object detection to evaluate the ability of open-vocabulary detectors to discern extrinsic object properties. Similarly, OVDEval (Yao et al. 2024) introduces an open-vocabulary detection benchmark to evaluate the performance on linguistic aspects using complex language prompts. Our work is different from them (Ramanathan et al. 2023; Bianchi et al. 2024; Yao et al. 2024) since we focus on the understanding of object attributes on 3D data, which poses greater challenges compared to the understanding in 2D images due to the limited annotations in 3D benchmarks.
F Limitations and Future Work
Our benchmark is currently constructed solely on the ScanNet200 benchmark with limited 3D indoor scene. It would be beneficial to increase the scale of our benchmark to include a wider variety of 3D scenes and objects. In future work, we plan to extend our OpenScan benchmark to encompass more diverse scenes by incorporating indoor 3D datasets such as ScanNet++ (Yeshwanth et al. 2023) and Matterport3D (Chang et al. 2017). Our mature annotation procedures can be readily adapted to these datasets. Moreover, we aim to evaluate current OV-3D models on our GOV-3D task, particularly examining performance variations when using higher point resolutions in ScanNet++ (Yeshwanth et al. 2023) and larger scene areas in Matterport3D (Chang et al. 2017).
G Broader Impact